1. In the laboratory a student finds that it takes 81.6 Joules to increase the temperature of 14.1 grams of solid iodine from 25.0 to 39.4 degrees Celsius.

The specific heat of iodine calculated from her data is  J/g°C.

2. In the laboratory a student finds that it takes 116 Joules to increase the temperature of 11.8 grams of solid sulfur from 24.0 to 38.1 degrees Celsius.

The specific heat of sulfur calculated from her data is  J/g°C.

5.1
Answer #1

1)

Given:

Q = 81.6 J

m = 14.1 g

Ti = 25.0 oC

Tf = 39.4 oC

use:

Q = m*C*(Tf-Ti)

81.6 = 14.1*C*(39.4-25.0)

C = 0.402 J/g.oC

Answer: 0.402 J/g.oC

2)

Given:

Q = 116 J

m = 11.8 g

Ti = 24 oC

Tf = 38.1 oC

use:

Q = m*C*(Tf-Ti)

116.0 = 11.8*C*(38.1-24.0)

C = 0.697 J/g.oC

Answer: 0.697 J/g.oC